# F-Bench: Rethinking Human Preference Evaluation Metrics for Benchmarking Face Generation, Customization, and Restoration

### v1

@@ 6 Conclusion
Evaluating the quality of AI-generated face images is a critical area of research in the development of generative models, given the human tendency to focus on faces. In this work, we introduce FaceQ, a meticulously curated AI face dataset comprising 12255 face images generated by 29 popular generative models, along with 32742 human annotated MOS scores. This dataset serves as a benchmark for evaluating the capabilities of existing face generation, customization, and restoration models, while also assessing the performance of current Image Quality Assessment (IQA) and Face Quality Assessment (FQA) methods. Our analysis highlights the unique features, strengths, and weaknesses of current face-related generative models, and reveals the limitations of traditional IQA and FQA algorithms in delivering subjectively consistent quality assessments for AI-generated faces. We believe FaceQ will contribute to advancing generative models and help overcome current limitations in face generation.
@@
Our evaluation framework has some limitations. First, we are mindful of privacy concerns related to face datasets, and all data collection and sharing adhere to relevant privacy policies. Second, while the maximum resolution in our dataset is limited to 1024 pixels, we recognize that future work may incorporate higher-resolution images as generative models continue to advance. We hope that this work will contribute to improving the quality of AI-generated faces and foster the development of objective face evaluation metrics in generative modeling.
Appendix A FaceQ: Dataset Construction
A.1 Face Generation, Customization and Restoration Models
@@
Tab. A.1 provides a comprehensive summary of models evaluated for face generation, editing, and restoration, including the model links, released dates, the resolution, and the backbone architectures. (1) Face generation. All the generation models are inference by pre-trained checkpoints in their default resolutions and hyper-parameters. Specifically, Stable Diffusion V1.5 [68], DreamLike [13], and RealisticVision [65] support high-resolution generation, such as 1024 1024, but we utilize their default training resolution due to severe subject repetition phenomenon. Deep Floyd [11] is a 3-stage pixel space diffusion model, here we only consider the third-stage results. For the dynamic step sampling, the number of steps per stage for StableCascade [58] and Deep Floyd [11] is reduced to the quarter. (2) Face customization. All the customization models are inference by pre-trained checkpoints in their default resolutions and hyper-parameters. FastComposer [97], originally designed for multi-subject customization, is assessed here with a single reference image as input. We use the IP-Adapter release version, including IP-Adapter-FaceID-SDXL [29] and IP-Adapter-FaceID-PlusV2 [28] (referred to as IP-Adapter-FaceID and IP-Adapter-FaceID-Plus). Their backbones are SDXL and SD-v1.5, respectively. (3) Face restoration. For DR2 [94], we follow the hyperparameter settings recommended in the original paper: for real-world inputs and for synthetic inputs.
@@
| Face Generation | Stable Diffusion V1.5 [68] | 2022.04 | Latent Diffusion | |
| Stable Diffusion V2.1 [68] | 2022.12 | Latent Diffusion | ||
@@
The face-centric prompts used for face generation can be categorized into nine classes. Tab. 6 presents two example prompts for each category due to space constraints. We ensured equal numbers of prompts for male and female subjects. The prompts for face customization are a subset of those used for face generation.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Degradation Scheme.
We construct two synthetic degradation pipelines to mimic real-world degradation. The first is first order pipeline following previous works [4, 92, 94] which can be expressed as
@@
A.2 Additional Examples of FaceQ Database
@@
A.3 Quantitative Analysis of FaceQ Database
@@
Appendix B FaceQ: Subjective Experiments
B.1 Implementation Details
@@
B.2 Subjective Evaluation Examples
@@
Appendix C F-Bench: More Analysis
C.1 MOS Distribution
@@
C.2 Perspectiive Analysis
To provide a clearer comparison of the strengths and weaknesses of different methods, we present the average MOS scores across various dimensions in Figure 16. For the three dimensions of face generation, authenticity exhibits the largest disparity between methods, while correspondence and quality tend to cluster around higher scores. In the face customization task, the methods show inconsistent performance in correspondence, whereas quality remains relatively balanced. For face restoration, quality-synthetic emerges as the easiest metric to achieve high scores, followed by quality-real-world. Figure 17 displays the rankings of the various methods. For face generation, Flux [14] achieves the highest performance across all three dimensions. When considering authenticity, RealisticVision [65] and SD3 [72] outperform other methods. Playground [39] ranks second only to Flux [14] in terms of correspondence, while Kolors [36] and SD3 [72] follow Flux [14] in quality. On the other hand, SDv1.5 [67] performs the worst across all dimensions. For face customization, IP-Adapter-FaceID-Plus [28], InstantID [88], and PhotoMaker [74] excel at preserving identity information. For face restoration, CodeFormer [107] demonstrates the best performance in synthetic scenarios, while StableSR achieves the highest scores in real-world scenarios.
C.3 Class-wise Comparison
@@
Appendix D QA Methods Implementation Details
Evaluation Metrics
@@ Appendix D QA Methods Implementation Details
SRCC, which ranges from -1 to 1, evaluates the monotonic relationship between two variables. For images, it is computed as:
Here, represents the rank of the ground truth value , while corresponds to the rank of the predicted value . When the SRCC value is higher, it signifies a stronger monotonic agreement between the ground truth and the predicted scores. PLCC quantifies the linear correlation between predicted scores and ground truth scores and is formulated as:
where and denote the mean values of the ground truth scores and the predicted scores, respectively. KLCC measures the ordinal association between two measured quantities and is defined as:

### v2

@@ 7 Conclusion
In this work, we conduct a pioneering generative face image quality assessment study. Specifically, we introduce FaceQ, the first generated face quality assessment database comprising 12,255 face images and 491,130 human preference ratings. Then, we establish F-Bench, a benchmark for evaluating the capabilities of existing face generation, customization, and restoration models, while also reveals the limitations of current Image Quality Assessment (IQA) and Face Quality Assessment (FQA) methods. Furthermore, we propose an evaluation model F-Eval to accurately predict human preferences on generated face images, which demonstrates state-of-the-art performance. We believe this work will contribute to advancing generative models and help overcome current limitations in AIGC community.
Acknowledgment
This work is supported by the National Natural Science Foundation of China under Grants 62271308, 62401365, 62225112, 62132006, U24A20220, STCSM under Grants 24ZR1432000, 24511106902, 24511106900, 22511105700, 22DZ2229005, 111 plan under Grant Number BP0719010, China Postdoctoral Science Foundation under Grants BX20250411, 2025M773473, and State Key Laboratory of UHD Video and Audio Production and Presentation.
Supplementary Material
@@ Acknowledgment
First, we are mindful of privacy concerns related to face datasets, and all data collection and sharing adhere to relevant privacy policies. Second, while the maximum resolution in our dataset is limited to 1024 pixels, we recognize that future work may incorporate higher-resolution images as generative models continue to advance.
8 FaceQ: Dataset Construction
8.1 Face Generation, Customization and Restoration Models
@@
Tab. 8.1 provides a comprehensive summary of models evaluated for face generation, editing, and restoration, including the model links, released dates, the resolution, and the backbone architectures. (1) Face generation. All the generation models are inference by pre-trained checkpoints in their default resolutions and hyper-parameters. Specifically, Stable Diffusion V1.5 [74], DreamLike [2], and RealisticVision [8] support high-resolution generation, such as 1024 1024, but we utilize their default training resolution due to severe subject repetition phenomenon. Deep Floyd [1] is a 3-stage pixel space diffusion model, here we only consider the third-stage results. For the dynamic step sampling, the number of steps per stage for StableCascade [70] and Deep Floyd [1] is reduced to the quarter. Each model’s negative prompt is configured to exclude “anime” and “semi-realistic” outputs by using terms like “worst quality”, “low quality”, “illustration”, “3D”, “2D”, “painting”, “cartoons”, “sketch”, “anime”, “animation”, “cartoon”, and “semi-realistic”. Safe sensors are enabled to filter out NSFW content. (2) Face customization. All the customization models are inference by pre-trained checkpoints in their default resolutions and hyper-parameters. FastComposer [104], originally designed for multi-subject customization, is assessed here with a single reference image as input. We use the IP-Adapter release version, including IP-Adapter-FaceID-SDXL [4] and IP-Adapter-FaceID-PlusV2 [5] (referred to as IP-Adapter-FaceID and IP-Adapter-FaceID-Plus). Their backbones are SDXL and SD-v1.5, respectively. (3) Face restoration. For DR2 [100], we follow the hyperparameter settings recommended in the original paper: for real-world inputs and for synthetic inputs.
@@
| Generation | Stable Diffusion V1.5 [rombach2022highresolutionimagesynthesislatent] | 2022.04 | Latent Diffusion | |
| Stable Diffusion V2.1 [rombach2022highresolutionimagesynthesislatent] | 2022.12 | Latent Diffusion | ||
@@
| Clothing | a woman wearing a purple wizard outfit |
@@
| Expression | a man crying disappointedly, with tears flowing |
@@
| View | a man wearing a doctoral cap, upper body, with the left side of the face facing the camera |
@@
| Style | instagram photo, portrait photo of a man, perfect face, natural skin, hard shadows, film grain |
@@
| Facial Attributes | a young girl with large round blue eyes, a flat nose bridge, and purple lipstick |
@@
The face-centric prompts used for face generation can be categorized into nine classes. Tab. 7 presents two example prompts for each category due to space constraints. We ensured equal numbers of prompts for male and female subjects. The prompts for face customization are a subset of those used for face generation. Degradation Scheme. We construct two synthetic degradation pipelines to mimic real-world degradation. The first is first order pipeline following previous works [15, 98, 100] which can be expressed as
@@
8.2 Additional Examples of FaceQ Database
@@
8.3 Quantitative Analysis of FaceQ Database
@@
9 FaceQ: Subjective Experiments
9.1 Implementation Details
@@
9.2 Subjective Evaluation Examples
@@
10 F-Bench: More Analysis
10.1 MOS Distribution
@@
10.2 Perspectiive Analysis
To provide a clearer comparison of the strengths and weaknesses of different methods, we present the average MOS scores across various dimensions in Figure 17. For the three dimensions of face generation, authenticity exhibits the largest disparity between methods, while correspondence and quality tend to cluster around higher scores. In the face customization task, the methods show inconsistent performance in correspondence, whereas quality remains relatively balanced. For face restoration, quality-synthetic emerges as the easiest metric to achieve high scores, followed by quality-real-world. Figure 18 displays the rankings of the various methods. For face generation, Flux [3] achieves the highest performance across all three dimensions. When considering authenticity, RealisticVision [8] and SD3 [9] outperform other methods. Playground [52] ranks second only to Flux [3] in terms of correspondence, while Kolors [6] and SD3 [9] follow Flux [3] in quality. On the other hand, SDv1.5 [rombach2022high] performs the worst across all dimensions. For face customization, IP-Adapter-FaceID-Plus [5], InstantID [94], and PhotoMaker [79] excel at preserving identity information. For face restoration, CodeFormer [116] demonstrates the best performance in synthetic scenarios, while StableSR achieves the highest scores in real-world scenarios.
10.3 Class-wise Comparison
@@
11 QA Methods Implementation Details
11.1 Evaluation Metrics
@@
SRCC., which ranges from -1 to 1, evaluates the monotonic relationship between two variables. For images, it is computed as:
Here, represents the rank of the ground truth value , while corresponds to the rank of the predicted value . When the SRCC value is higher, it signifies a stronger monotonic agreement between the ground truth and the predicted scores. PLCC. quantifies the linear correlation between predicted scores and ground truth scores and is formulated as:
where and denote the mean values of the ground truth scores and the predicted scores, respectively. KLCC. measures the ordinal association between two measured quantities and is defined as:
@@
12 More Details of Our F-Eval Model
12.1 Loss Funtion
We use both language loss, L1 loss and cross-entropy loss as the loss functions to optimize the training process. Specifically, the language loss is used to restrict the F-Eval to produce specific quality-related answer patterns. The language loss function can be formulated as:
where is the predicted token, is the ground truth token, indicates the probability, and is the number of tokens. L1 loss is used to regress the quality scores. The L1 loss can be formulated as:
where is the predicted quality score, is the ground truth quality score, and is the number of images in a batch. Cross-entropy loss to predict the dimension ID from the input text tokens. The cross-entropy loss can be formulated as:
where is the dimension label for the -th sample in class (dimension ID), is the predicted probability for the -th sample in class , and is the number of images in a batch. is 4. The overall loss function can be formulated as:
| Task | Face Generation | Face Restoration syn | |||||||||||||
| Dimension | Quality | Authenticity | Correspondence | Quality | ID Fidelity | ||||||||||
| Method | SRCC | KRCC | PLCC | SRCC | KRCC | PLCC | SRCC | KRCC | PLCC | SRCC | KRCC | PLCC | SRCC | KRCC | PLCC |
| Freeze Projector | 0.8080 | 0.6428 | 0.8698 | 0.7518 | 0.5832 | 0.8123 | 0.8475 | 0.6783 | 0.9031 | 0.7769 | 0.5787 | 0.8074 | 0.7757 | 0.5832 | 0.8074 |
| Freeze Face encoder | 0.8151 | 0.6450 | 0.8733 | 0.7524 | 0.5834 | 0.8140 | 0.8347 | 0.6654 | 0.8912 | 0.8121 | 0.6120 | 0.8430 | 0.7828 | 0.6021 | 0.8341 |
| Single LoRA | 0.7927 | 0.6140 | 0.8659 | 0.7757 | 0.5995 | 0.8285 | 0.8191 | 0.6429 | 0.8760 | 0.8377 | 0.6538 | 0.8874 | 0.8228 | 0.6341 | 0.8556 |
| Task LoRA | 0.8231 | 0.6477 | 0.8895 | 0.7803 | 0.6056 | 0.8317 | 0.8377 | 0.6661 | 0.8943 | 0.8244 | 0.6428 | 0.8806 | 0.8257 | 0.6359 | 0.8519 |
| F-Eval (Ours) | 0.8486 | 0.6670 | 0.9085 | 0.8312 | 0.6585 | 0.8578 | 0.8471 | 0.6637 | 0.9106 | 0.8692 | 0.6855 | 0.9009 | 0.8507 | 0.6731 | 0.8726 |
| Task | Face Customization | Face Restoration rw | |||||||||||||
| Dimension | Quality | Authenticity | Correspondence | Quality | ID Fidelity | ||||||||||
| Method | SRCC | KRCC | PLCC | SRCC | KRCC | PLCC | SRCC | KRCC | PLCC | SRCC | KRCC | PLCC | SRCC | KRCC | PLCC |
| Freeze Projector | 0.9242 | 0.7651 | 0.9233 | 0.9269 | 0.7679 | 0.9242 | 0.8429 | 0.6726 | 0.8386 | 0.7067 | 0.5249 | 0.7114 | 0.4778 | 0.3616 | 0.5451 |
| Freeze Face encoder | 0.9239 | 0.7662 | 0.9217 | 0.9214 | 0.7612 | 0.9208 | 0.8550 | 0.6836 | 0.8524 | 0.7158 | 0.5421 | 0.8532 | 0.6028 | 0.4386 | 0.6898 |
| Single LoRA | 0.9419 | 0.7938 | 0.9421 | 0.9424 | 0.7957 | 0.9430 | 0.8850 | 0.7140 | 0.8974 | 0.8019 | 0.6155 | 0.8455 | 0.6710 | 0.5017 | 0.7768 |
| Task LoRA | 0.9421 | 0.7984 | 0.9403 | 0.9422 | 0.7985 | 0.9402 | 0.8727 | 0.7005 | 0.8918 | 0.7540 | 0.5661 | 0.7964 | 0.6376 | 0.4736 | 0.7403 |
| F-Eval (Ours) | 0.9462 | 0.7961 | 0.9461 | 0.9188 | 0.7640 | 0.9322 | 0.9460 | 0.7959 | 0.9457 | 0.8448 | 0.6577 | 0.8705 | 0.7957 | 0.6057 | 0.8366 |
12.2 Failure Cases
F-Eval’s performance may degrade when the face region in the image is either too small or in an extreme side view. In these scenarios, the face features extracted by the face encoder become less accurate, leading to a drop in performance, particularly in the quality and authenticity dimensions. We will address this limitation by integrating a more robust face encoder fine-tuned on side view and small face data in future works. Additionally, F-Eval currently does not support identifying specific distorted regions, which will also be addressed in future works.
12.3 More Experimental Results
The detailed ablation study results are listed in Tab. 8. The results indicate the effectiveness of the key components in our F-Eval.